Tohban Report 2016-03-02
Jump to navigation
Jump to search
Tohban Reports | |
---|---|
Start Date: | 24 Feb 2016 |
End Date: | 2 Mar 2016 |
Tohban: | Hazel Bain |
Tohban email: | hbain@ssl.berkeley.edu |
Next Tohban: | Martin Fivian |
List all reports |
Solar Activity
Solar activity this week was low.
How many GOES flares occurred?
Flares above B, C, M, X class were 27 3 0 0
RHESSI flare list contains 0 over the time range 24-Feb-16 02-Mar-16
Memory Management
SSR is currently emptying at the end of the pass set.
Spacecraft Status
The spacecraft is currently in the middle of an anneal procedure.
Data Gaps
N/A
Detector issues
The detectors are currently off.
Spacecraft Management
Decimation | N/A |
HLAT Decimation | N/A |
Night time data (fronts) | N/A |
Night time data (rears) | N/A |
Require extra passes? | N/A |
Requirement for moving pointer? | N/A |
Attenuator operation | N/A |
Detector problems? | N/A |